an alpha particle is projected from infinity with Vo towards the nucleus of gold atom... find the velocity of alpha particle at a distended of 2Ro....

Grade:11

4 Answers

Anish
2 years ago
Kinetic energy = potential energy
½*m*vo2 = k*2e*Ze/ro – (1)
and let velocity is v
then
½*m*v2 = k*2e*Ze/2ro -(2)
then (2)/(1)
v2/vo2 = ½
then v = vo/\sqrt{2}
Hope it clears.
 
Anish
2 years ago
Kinetic energy = potential energy
½*m*vo2 = k*2e*Ze/ro – (1)
and let velocity is v
then
½*m*v2 = k*2e*Ze/2ro -(2)
then (2)/(1)
v2/vo2 = ½
then v = vo/\sqrt{2}
Hope it is clear :) 
Anish
2 years ago
Kinetic energy = potential energy
½*m*vo2 = k*2e*Ze/ro – (1)
and let velocity is v
then
½*m*v2 = k*2e*Ze/2ro -(2)
then (2)/(1)
v2/vo2 = ½
then v = vo/\sqrt{2}
Hope it is clear. 
Anish
2 years ago
Kinetic energy = potential energy
½*m*vo2 = k*2e*Ze/ro – (1)
and let velocity is v
then
½*m*v2 = k*2e*Ze/2ro -(2)
then (2)/(1)
v2/vo2 = ½
then v = vo/\sqrt{2}
Hope it is clear